OpenStudy (anonymous):

Find the volume of half a sphere with the radius of 10.

OpenStudy (abb0t):

Do you know the formula for volume of a sphere? It's \(\large V= \frac{4}{3}\pi r^3\) where: r = radius Now, for the volume of the half sphere, do as it says, cut it in half, which means, divide the volume of a sphere by 2.

OpenStudy (anonymous):

So what would you get as the answer?

OpenStudy (abb0t):

:) What did \(\huge YOU\) get?

OpenStudy (anonymous):

something in the 2000's. which i know is wrong

OpenStudy (anonymous):

isn't just 3.14 times radius^3 divide by two

OpenStudy (abb0t):

Yes, you're right. I got: \(2094.4\)

OpenStudy (abb0t):

As long as you plug everything in correctly into your calculator, you're answer should be correct. You're simply plugging your given value of \(10^3\) which is same thing as 1000

OpenStudy (anonymous):

Okay well I moved onto a different problem. Using the quadratic formula, solve x^2+5x-14=0. & approximate to the nearest tenth. I got -4,7 but it says Im wrong. Am I?

OpenStudy (abb0t):

Yes, you are wrong. One of your values IS close but is actually supposed to be -7. Anyways, here is the quadratic formula for when you have: \(ax^2+bx+c\) it is: \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (abb0t):

I don't know if it's displaying properly on your screen, but the root should be all over like this: \(\sqrt{b^2-4ac}\)

OpenStudy (abb0t):

If you plug in everything: \[\frac{ -5 \pm \sqrt{(5)^2 - 4(-14)(1)} }{ 2(1) }\]

OpenStudy (dan815):

abbot is that volume formula right for spherical coords?

OpenStudy (anonymous):

Thats how I have it.

OpenStudy (anonymous):

I give up

OpenStudy (dan815):

why wud u give up >_> u just need to volume of a spher with radius 10 and divide by 2 -.-

OpenStudy (abb0t):

@dan815 please do not spam in other people questions. And yes, in sperical coordinates, I think that's how you would prove the volume of a sphere..but that's irrelevant here obviously.

OpenStudy (abb0t):

@amanda_deason you need to plug in the given values into your calculator, you should get: \[\frac{ -5 \pm \sqrt{81} }{ 2 }\]

OpenStudy (anonymous):

@abb0t it's just frustrating because this program keeps saying im getting it wrong when its actually right. i just wish that i could have all the answers because im supposed to graduate tomorrow and now i cant.

OpenStudy (dan815):

do u wanna learn complete the square? thats where quadratic formula comes from

OpenStudy (abb0t):

I think you forgot to divide by \(\huge 2\) because if you notice, \(-5+9 = 4\) for the top and \(-5 - 9 = -14\) You SHOULD get: x = 2 and x = -7

OpenStudy (anonymous):

thats what i put and i still got it wrong.

OpenStudy (abb0t):

You put \(\large x = 2\) and \(\large x = -7\) and it says you're wrong?

OpenStudy (dan815):

then put X=-7 and X = 2 :)

OpenStudy (abb0t):

If that is the case, you need to discuss that with your school because those are the solutions to your quadratic function.

OpenStudy (anonymous):

okay thank you.